OpenStudy (anonymous):

the sum of two numbers is 36 the difference is 6 find the numbers

OpenStudy (aroub):

x+y=36 x-y=6 -------- 2x=60 x=30 x+y=36 30+y=36 y=6 The two numbers are 6 and 30.